Find the following and identify the domain:
given f(x)= x - 2 and g(x)=x(squared)- 4
solve for (f=g)(x)=

Answer by Nate(3500)

f(x) = x - 2 and g(x) = x^2 - 4
(f o g)(x) = x^2 - 4 - 2 = x^2 - 6